1. Using average bond enthalpies (linked above), estimate the enthalpy change for the following reaction:

2CO(g) + 2NO(g)2CO2(g) + N2(g)

___ kJ

2. Using average bond enthalpies (linked above), estimate the enthalpy change for the following reaction:

H2(g) + I2(g)2HI(g)

___ kJ
1
Expert's answer
2019-11-25T07:25:08-0500

Solution.

"\Delta H = \Sigma(\Delta H(broken \ bonds)) - \Sigma(\Delta H(formed \ bonds))"

1.

Divide the coefficients in the reaction equation:

"CO + NO = CO2 + \frac{N2}{2}"

"\Delta H = 1072 + 628 - (799 \times 2) - \frac{1 \times 941}{2} = -368.5 \ kJ/mol"

2.

This equation can be left the same.

"\Delta H = 436 + 151 - (2 \times 299) = -11 \ kJ/mol"

Answer:

1.

"\Delta H = -368.5 \ kJ/mol"

2.

"\Delta H = -11 \ kJ/mol"